Hi!
How would I solve the square root of a number when there's another number on the outside of the square root? For example, the example is asking me to find the decimal expression of the square root of 64, which I know is 8, but then it has the number 4 on the outside of the square root? I'm not sure what to do with the 4. Thank you!

Answer by stanbon(75887)

How would I solve the square root of a number when there's another number on the outside of the square root? For example, the example is asking me to find the decimal expression of the square root of 64, which I know is 8, but then it has the number 4 on the outside of the square root? I'm not sure what to do with the 4. Thank you!
--------------------
4*sqrt(64) = 4*8 = 32
---
If the "4" is small and indicates "take the 4th root",
the answer would be:
4th root of (64) = 4th root of 2^6 = 2^(6/4) = 2^(3/2)
==========================================================
Cheers,
Stan H.
===========================